# An Extensible Framework for Open Heterogeneous Collaborative Perception

### v1

@@ Abstract
Collaborative perception aims to mitigate the limitations of single-agent perception, such as occlusions, by facilitating data exchange among multiple agents. However, most current works consider a homogeneous scenario where all agents use identity sensors and perception models. In reality, heterogeneous agent types may continually emerge and inevitably face a domain gap when collaborating with existing agents. In this paper, we introduce a new open heterogeneous problem: how to accommodate continually emerging new heterogeneous agent types into collaborative perception, while ensuring high perception performance and low integration cost? To address this problem, we propose HEterogeneous ALliance (HEAL), a novel extensible collaborative perception framework. HEAL first establishes a unified feature space with initial agents via a novel multi-scale foreground-aware Pyramid Fusion network. When heterogeneous new agents emerge with previously unseen modalities or models, we align them to the established unified space with an innovative backward alignment. This step only involves individual training on the new agent type, thus presenting extremely low training costs and high extensibility. It also protects new agents’ model details from disclosure since the training can be conducted by the agent owner locally. To enrich agents’ data heterogeneity, we bring OPV2V-H, a new large-scale dataset with more diverse sensor types. Extensive experiments on OPV2V-H and DAIR-V2X datasets show that HEAL surpasses SOTA methods in performance while reducing the training parameters by 91.5% when integrating 3 new agent types. Code and data are available at: https://github.com/yifanlu0227/HEAL.
@@ 1 Introduction
  • To train the collaboration base, we introduce a powerful Pyramid Fusion. It utilizes multiscale and foreground-aware designs to realize robust multi-agent feature fusion and sustains a potent unified feature space for all agents.
  • To integrate new types of agents, we propose a novel backward alignment mechanism to align heterogeneous agents to the unified space. This training is conducted locally on single agents, reducing training costs while also preserving model details.
@@ 4.2 New Agent Type Training
Specifically, we denote the new agent type as $n_{1}$ and new agent set as $\mathcal{S}_{[n_{1}]}$, and the full set of current agents set becomes $\mathcal{S}=\mathcal{S}_{[b]}\cup\mathcal{S}_{[n_{1}]}$. For agent $k$ in the agent set $\mathcal{S}_{[n_{1}]}$, we define $\mathbf{O}_{k}$ as its observation and $f_{\mathrm{encoder}[n_{1}]}(\cdot)$ as its detector encoder. We keep $f^{*}_{\rm pyramid\_fusion}(\cdot)$ and $f^{*}_{\rm head}(\cdot)$ from previous stage unchanged, where *** denotes fixed, and train $f_{\mathrm{encoder}[n_{1}]}(\cdot)$ on single agents:
@@ 5.1 Datasets
OPV2V-H. We propose a simulation dataset dubbed OPV2V-H. In OPV2V (Xu et al., 2022c) dataset, the LiDAR with 64-channel shows a significant detection advantage over the camera modality. Evaluations of heterogeneous collaboration on OPV2V may not truly represent the collaboration performance between these two modalities since LiDAR agents provide most detections (Xiang et al., 2023). For this purpose, we collected more data to bridge the gap between LiDAR and camera modalities, leading to the new OPV2V-H dataset. We utilize CARLA (Dosovitskiy et al., 2017) and OpenCDA (Xu et al., 2021) for our dataset simulation.
OPV2V-H dataset has on average approximately 3 agents with a minimum of 2 and a maximum of 7 in each frame. Except for one 64-channel LiDAR and four 4 cameras of resolution 800*600 of each agent from original OPV2V, OPV2V-H collects 16-channel and 32-channel LiDAR data as well as 4 depth camera data. This ensures a more balanced data quality between LiDAR and cameras, and having ground-truth depth is highly beneficial for camera-based detection learning.
@@ 5.2 Open Heterogeneous settings
[…]
@@
[…]
@@
[…]
@@ 5.3 Quantitative Results
Figure 4: Robust Experiment to pose error and compression ratio. Pose noise is set to $\mathcal{N}(0,\sigma_{p}^{2})$ on x, y position and $\mathcal{N}(0,\sigma_{r}^{2})$ on yaw angle. HEAL achieves SOTA under various degraded conditions.
@@ 6 Conclusion
This paper proposes HEAL, a novel framework for open heterogeneous collaborative perception. HEAL boasts exceptional collaborative performance, minimal training costs and model-detail protection. Experiments on our proposed OPV2V-H and DAIR-V2X datasets validate the efficiency of HEAL, offering a practical solution for extensible collaborative perception deployment in the real-world scenario. The limitation of HEAL is that it requires BEV features, which requires extra effort for some model (e.g., keypoint-based LiDAR detection) to be compatible to the framework.

### v3

@@ Abstract
Collaborative perception aims to mitigate the limitations of single-agent perception, such as occlusions, by facilitating data exchange among multiple agents. However, most current works consider a homogeneous scenario where all agents use identity sensors and perception models. In reality, heterogeneous agent types may continually emerge and inevitably face a domain gap when collaborating with existing agents. In this paper, we introduce a new open heterogeneous problem: how to accommodate continually emerging new heterogeneous agent types into collaborative perception, while ensuring high perception performance and low integration cost? To address this problem, we propose HEterogeneous ALliance (HEAL), a novel extensible collaborative perception framework. HEAL first establishes a unified feature space with initial agents via a novel multi-scale foreground-aware Pyramid Fusion network. When heterogeneous new agents emerge with previously unseen modalities or models, we align them to the established unified space with an innovative backward alignment. This step only involves individual training on the new agent type, thus presenting extremely low training costs and high extensibility. To enrich agents’ data heterogeneity, we bring OPV2V-H, a new large-scale dataset with more diverse sensor types. Extensive experiments on OPV2V-H and DAIR-V2X datasets show that HEAL surpasses SOTA methods in performance while reducing the training parameters by 91.5% when integrating 3 new agent types. We further implement a comprehensive codebase at: https://github.com/yifanlu0227/HEAL.
@@ 1 Introduction
  • We propose a powerful Pyramid Fusion for the collaboration base training, which utilized multiscale and foreground-aware designs to sustain a potent unified feature space.
  • To integrate new types of agents, we introduce a novel backward alignment mechanism to align heterogeneous agents to the unified space. This training is conducted locally on single agents, reducing training costs while also preserving model details.
@@ 4.2 New Agent Type Training
Specifically, we denote the new agent type as $n_{1}$ and new agent set as $\mathcal{S}_{[n_{1}]}$, and the full set of current agents set becomes $\mathcal{S}=\mathcal{S}_{[b]}\cup\mathcal{S}_{[n_{1}]}$. For agent $k$ in the agent set $\mathcal{S}_{[n_{1}]}$, we define $\mathbf{O}_{k}$ as its observation and $f_{\mathrm{encoder}[n_{1}]}(\cdot)$ as its detector encoder. We keep $f^{*}_{\rm pyramid\_fusion}(\cdot)$ and $f^{*}_{\rm head}(\cdot)$ from previous stage unchanged, where * denotes fixed, and train $f_{\mathrm{encoder}[n_{1}]}(\cdot)$ on single agents:
@@ 5.1 Datasets
OPV2V-H. We propose a simulation dataset dubbed OPV2V-H. In OPV2V (Xu et al., 2022c) dataset, the LiDAR with 64-channel shows a significant detection advantage over the camera modality. Evaluations of heterogeneous collaboration on OPV2V may not truly represent the collaboration performance between these two modalities since LiDAR agents can provide most detections (Xiang et al., 2023). For this purpose, we collected more data to bridge the gap between LiDAR and camera modalities, leading to the new OPV2V-H dataset. OPV2V-H dataset has on average approximately 3 agents with a minimum of 2 and a maximum of 7 in each frame. Except for one 64-channel LiDAR and four 4 RGB cameras (resolution 800*600) of each agent from original OPV2V dataset, OPV2V-H collects extra 16- and 32-channel LiDAR data and 4 depth camera data.
@@ 5.2 Open Heterogeneous settings
Metric AP50 $\uparrow$ AP70 $\uparrow$ Train Throughput (#/sec.) $\uparrow$
@@
Metric Model #Params (M) $\downarrow$ FLOPs(T) $\downarrow$ Peak Memory (GB) $\downarrow$
@@
Metric AP70 $\uparrow$ $\sum$ M.#P. $\downarrow$ AP70 $\uparrow$ $\sum$ M.#P. $\downarrow$ AP50 $\uparrow$ $\sum$ M.#P. $\downarrow$ AP50 $\uparrow$ $\sum$ M.#P. $\downarrow$ AP50 $\uparrow$ $\sum$ M.#P. $\downarrow$
@@ 5.3 Quantitative Results
Figure 4: Robust Experiment to pose error and compression ratio. Pose noise is set to $\mathcal{N}(0,\sigma_{p}^{2})$ on x, y location and $\mathcal{N}(0,\sigma_{r}^{2})$ on yaw angle.
@@ 6 Conclusion
This paper proposes HEAL, a novel framework for open heterogeneous collaborative perception. HEAL boasts exceptional collaborative performance, minimal training costs and model-detail protection. Experiments on our proposed OPV2V-H and DAIR-V2X datasets validate the efficiency of HEAL, offering a practical solution for extensible collaborative perception deployment in the real-world scenario. The limitation of HEAL is that it requires BEV features, which requires extra effort for some models (e.g., keypoint-based LiDAR detection) to be compatible to the framework.
7 Acknowledge
This research is supported by the National Key R&D Program of China under Grant 2021ZD0112801, NSFC under Grant 62171276 and the Science and Technology Commission of Shanghai Municipal under Grant 21511100900, 22511106101 and 22DZ2229005.